OpenStudy (anonymous):

help please ! 5+3/|2-3x| < 8

hero (hero):

\[5 + \frac{3}{|2 - 3x|} < 8\]

hero (hero):

Is that what you're working on?

hero (hero):

So anyway, I'm just going to post the solution assuming that what I have posted is correct .

OpenStudy (anonymous):

yup , sorry for late reply .

hero (hero):

\[ \frac{3}{|2 - 3x|} < 8 - 5 \\\frac{3}{|2 - 3x|} < 3 \\\frac{1}{|2 - 3x|} < \frac{3}{3} \\\frac{1}{|2 - 3x|} < 1 \\1 < |2 - 3x| \\|2 - 3x| > 1 \\|3x - 2| > 1 \\3x - 2 > \pm 1 \\-1> 3x - 2 > 1 \\-1 + 2 > 3x > 1 + 2 \\1 > 3x > 3 \\1/3 > x > 3/3 \\1/3 > x > 1 \\ x > 1 \text{ or } x < 1/3 \]

OpenStudy (anonymous):

okay thank you for the help . Really appreciate it

hero (hero):

I hope you understood the steps. You can do it several ways.

hero (hero):

There are a couple of rules of use to know like for example |x - y| = |y - x|

hero (hero):

for example if x = 5 and y = 9, then |5 - 9| = |-4| = 4 |9 - 5| = | 4| = 4

hero (hero):

I like that rule because I never have to worry about dividing by negative numbers. I never divide by a negative number.

OpenStudy (anonymous):

okay , but how about this question .

OpenStudy (anonymous):

1/( x+2 ) > - 1/( x -2)

OpenStudy (anonymous):

umm , can you show the calculations ?

hero (hero):

\[\frac{1}{x+2} > -\frac{1}{x-2} \\\space{}\\ \\\text{Multiply the fraction on the left side by (x+2)/(x+2): }\\\space{}\\ \\\frac{x+2}{(x+2)^2} > -\frac{1}{x-2} \\\space{}\\ \text{Cross Multiply:} \\\space{}\\ \\(x+2)(x-2) > -(x+2)^2 \\\space{}\\ \text{Expand:} \\\space{}\\ \\x^2 - 4 >-(x^2 + 4x + 4) \\x^2 - 4 > -x^2 - 4x - 4 \\x^2 + x^2+4x + 4 - 4 > 0 \\2x^2 + 4x > 0 \\2x(x+2) > 0 \\2x > 0 \\x + 2 > 0 \] Obviously when we solve 2x > 0, x = 0, so 0 > 0 makes no sense, however we can use it as a critical point. Solving x + 2 >0 x > -2 So we have critical points -2, 0, and 2 Thus our interval is \[(-2,0), (2,\infty)\]